OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 4x – y = 4 x + 4y = 18

OpenStudy (anonymous):

(1, 4) / (4, 2) / (4, 1) / (2, 4)

myininaya (myininaya):

solve the first for either y or x say we solve for y now plug that y into the second equation where there is y and solve that equation for x and then go back to the first equation that you solved for y and plug the x you just found wherever there is an x

myininaya (myininaya):

\[-y=4-4x => y=-4+4x\] see i solved the first equation for y

myininaya (myininaya):

now plug this in wherever you have a y in the equation and solve for x

myininaya (myininaya):

\[x+4(-4+4x)=18\] solve this for x

myininaya (myininaya):

You know it would be best not to post the same question twice. Thanks.

OpenStudy (anonymous):

after solving for x, we get x=2